# TraCS: Trajectory Collection in Continuous Space under Local Differential Privacy

### v1

@@
Abstract.
@@
Trajectory collection is fundamental for location-based services but often involves sensitive information, such as a user’s daily routine, raising privacy concerns. Local differential privacy (LDP) provides provable privacy guarantees for users, even when the data collector is untrusted. Existing trajectory collection methods ensure LDP only for discrete location spaces, where the number of locations affects their privacy guarantees and trajectory utility. Moreover, the location space is often naturally continuous, such as in flying and sailing trajectories, making these methods unsuitable. This paper proposes two trajectory collection methods that ensure LDP for continuous spaces: TraCS-D, which perturbs the direction and distance of locations, and TraCS-C, which perturbs the Cartesian coordinates of locations. Both methods are theoretically and experimentally analyzed for trajectory utility. TraCS can also be applied to discrete spaces by rounding perturbed locations to the nearest discrete points. It is independent of the number of locations and has only $\Theta(1)$ time complexity in each perturbation generation. Evaluation results on discrete location spaces validate this advantage and show that TraCS outperforms state-of-the-art methods with improved trajectory utility, especially for large privacy parameters.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/ZhengYeah/TraCS.
1. Introduction
@@
A trajectory is a sequence of locations that describes movement over time. Users’ trajectory data serves as a fundamental data source for location-based applications, such as navigation and route planning. However, the direct collection of trajectory data from users raises significant privacy concerns, particularly given the growing privacy expectations of users and stringent regulations (gdp, 2016; Krumm, 2009). Anonymization techniques, such as $k$-anonymity (Sweeney, 2002; Li et al., 2007), have been shown to be insufficient for protecting trajectory privacy (Chen et al., 2011). Therefore, stronger privacy guarantees are needed for such data.
@@
Local differential privacy (LDP) provides a provable worst-case privacy guarantee by requiring the hardness of distinguishing the presence of a user’s data from any released data. The privacy level of an LDP mechanism is quantified by its privacy parameter $\varepsilon$, which also implies the extent of data perturbation. Through LDP, users can release their perturbed trajectory data to an untrusted data collector while still benefiting from the location-based services provided by the collector.
Existing LDP methods for trajectory collection are designed for discrete spaces. They either rely on properly splitting the entire location space into grids (Du et al., 2023) or assume the location space consists of a set of labeled locations (a.k.a. points of interest) (Cunningham et al., 2021b; Zhang et al., 2023). These methods’ privacy guarantees are w.r.t. their corresponding discretized sets, and their efficacy depends on the size of these sets. Improper discretization leads to a loss of privacy guarantees and reduced trajectory utility. Moreover, labeled locations are not always available, especially in continuous spaces such as those encountered in flying and sailing trajectories. Therefore, it is necessary to design effective trajectory collection methods for continuous spaces.
However, designing such methods is challenging due to the following reasons: (i) they must guarantee LDP w.r.t. each location in the continuous space; (ii) despite the rigorous privacy guarantee, they must still provide sufficient trajectory utility.
@@
This paper proposes two LDP methods for trajectory collection in continuous spaces. The main insight is to decompose the $2$D continuous space into $1$D subspaces and design mechanisms for each subspace. Our mechanisms leverage the concept of piecewise-based mechanisms (Wang et al., 2019; Li et al., 2020) for $1$D bounded numerical domains.  Based on two different decompositions of the continuous space, we propose two methods, TraCS-D and TraCS-C. TraCS-D treats the continuous space as the composition of a direction space and a distance space. We design a novel direction perturbation mechanism to ensure LDP for the direction space. Combined with a redesigned piecewise-based mechanism on the distance space, TraCS-D ensures LDP for the continuous $2$D space. TraCS-C treats the continuous space as the Cartesian product of two distance spaces. Each location has a $2$D Cartesian coordinate representation, and TraCS-C perturbs these coordinates.  Our methods can also be applied to discrete spaces by rounding the perturbed location to the nearest discrete points. Compared with existing methods based on the Exponential mechanism (Cunningham et al., 2021b; Zhang et al., 2023), our methods are not affected by the number of locations in the location space and cost only $\Theta(1)$ time in each perturbation generation. Specifically, our contributions are as follows:
@@
  • To the best of our knowledge, this is the first work to provide trajectory collection methods for continuous space under LDP. We propose TraCS-D and TraCS-C for this purpose.  These methods introduce novel location perturbation mechanisms that exploit the direction information and coordinate information of trajectories in continuous spaces. We theoretically and experimentally analyze their trajectory utility.
@@
  • Our methods TraCS can also be applied to discrete spaces. Compared with existing methods for discrete spaces, TraCS have significantly lower computation complexity in generating perturbed locations. Experimental results on the discrete location spaces validate the efficiency of TraCS, and also show that it outperforms the state-of-the-art methods with enhanced trajectory utility, particularly for large privacy parameters.
2. Preliminaries
@@
This section formulates the problem, introduces LDP, and reviews existing trajectory collection methods.
2.1. Problem Formulation
@@
A typical trajectory collection schema consists of a set of trajectories from users and one collector. Each trajectory $\mathcal{T}$ is a sequence of locations $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$, where $\tau_{i}\in\mathcal{S}$ and $\mathcal{S}\subset\mathbb{R}^{2}$ is a continuous and bounded domain. The collector needs to collect these trajectories to provide location-based services.
@@
However, the collector is untrusted and may act as an adversary attempting to infer users’ private data. Therefore, releasing the private trajectories to the collector pose a privacy risk. To protect privacy, users perturb their trajectories using a privacy mechanism $\mathcal{M}:\mathcal{S}\to Range(\mathcal{M})$, and then send the perturbed trajectories $\mathcal{T}^{\prime}=\{\tau_{1}^{\prime},\tau_{2}^{\prime},\dots,\tau_{n}^{\prime}\}$ to the collector.
@@
We aim to design a mechanism $\mathcal{M}$ that satisfies $\varepsilon$-LDP, providing a provable privacy guarantee for the private trajectories. Additionally, $\mathcal{M}$ should ensure the utility of the perturbed trajectories and incur limited computational cost for real-time applications.
2.2. Local Differential Privacy
Definition 2.1 ($\varepsilon$-LDP (Duchi et al., 2013)).
@@ 2.2. Local Differential Privacy
A perturbation mechanism $\mathcal{M}:\mathcal{X}\to Range(\mathcal{M})$ satisfies $\varepsilon$-LDP, iff for two arbitrary input $x_{1}$ and $x_{2}$, the probability ratio of outputting the same $y$ is bounded:
@@
If $\mathcal{M}(x)$ is continuous, the probability $\Pr[\cdot]$ is replaced by probability density function ($pdf$). Intuitively, Definition 2.1 represents the difficulty of distinguishing $x_{1}$ and $x_{2}$ given $y$. Lower values of the privacy parameter $\varepsilon\in[0,+\infty)$ mean higher privacy. For example, $\varepsilon=0$ requires $\mathcal{M}$ to map two arbitrary inputs to any output $y$ with the same probability, thus the perturbed output contains no distribution information of the private input, making any hypothesis-testing method to infer the private $x$ powerless.
Theorem 2.2 (Sequential Composition (Dwork and Roth, 2014; Zhang et al., 2023)).
@@
Sequential composition allows us to design LDP mechanisms for $2$D data by composing two mechanisms for $1$D.
2.2.1. Piecewise-based Mechanism
If $\mathcal{X}$ is continuous and bounded, the state-of-the-art LDP mechanisms are piecewise-based (Wang et al., 2019; Li et al., 2020), i.e. they sample the output from piecewise probability distributions defined on $\mathcal{X}$. We summarize them in the following form:
Definition 2.3 (Piecewise-based mechanism).
@@ 2.2.1. Piecewise-based Mechanism
where $p_{\varepsilon}$ is the sampling probability w.r.t. $\varepsilon$ and $[l_{x,\varepsilon},r_{x,\varepsilon})$ is the sampling interval w.r.t. $x$ and $\varepsilon$. $Range(\mathcal{M})\supseteq\mathcal{X}$ is also a continuous and bounded domain.
Piecewise-based mechanisms sample the output $y$ with a high probability $p_{\varepsilon}$ within $[l_{x,\varepsilon},r_{x,\varepsilon})$ and a lower probability in the remaining two pieces, while satisfying the $\varepsilon$-LDP constraint. Typical instantiations include PM (Wang et al., 2019) for mean estimation and SW (Li et al., 2020) for distribution estimation. Our perturbation mechanisms also leverage the piecewise-based design but with different instantiations tailed for different data domains.
2.2.2. $k$-RR
If $\mathcal{X}$ is discrete, especially when $|\mathcal{X}|$ is not large (Wang et al., 2017), a classical LDP mechanism is $k$-RR (Kairouz et al., 2014).
Definition 2.4.
$k$-RR is a sampling mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{X}$ that, given input $x\in\mathcal{X}$, outputs $y\in\mathcal{X}$ according to
@@ 2.2.2. $k$-RR
$k$-RR outputs the truth $x$ with a higher probability or outputs other values with a lower probability, while satisfying $\varepsilon$-LDP.
2.3. Existing Methods
Table 1. Comparison with existing methods.
@@ 2.3. Existing Methods
NGram (Cunningham et al., 2021b) Hierarchical decomposition Discrete space
LDPTrace (Du et al., 2023) Model-based synthesis
ATP (Zhang et al., 2023) Direction perturbation
This paper (TraCS) Direction* & coordinate perturbation Continuous space
@@
Table 1 summarizes existing LDP methods for trajectory collection and our methods.
[…]
LDPTrace (Du et al., 2023) synthesizes trajectories based on a generative model. Specifically, the physical space $\mathcal{S}$ is discretized into cells, and each private trajectory is represented as a sequence of cells $\mathcal{T}=\{c_{1},c_{2},\dots,c_{n}\}$. Neighboring cells, i.e. $c_{i}$ and $c_{i+1}$, form transition patterns, which are perturbed by LDP mechanisms. These perturbed transition patterns are then used to train a Markov chain model that captures the transition patterns between cells. Finally, new trajectories $\mathcal{T}^{\prime}$ are synthesized by the trained model.
ATP (Zhang et al., 2023) uses direction perturbation to limit the set of possible locations. It assumes the location space is a finite set of labeled locations and represents a trajectory $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$ as a sequence of these locations. When perturbing a location $\tau_{i+1}$, ATP perturbs the direction from $\tau_{i}$ to $\tau_{i+1}$. It divides the direction space into $k$ sectors and uses $k$-RR to sample a direction sector. The perturbed location $\tau_{i+1}^{\prime}$ should be within the sampled sector, thereby limiting the possible $\tau_{i+1}^{\prime}$, which is then sampled using the Exponential mechanism.
@@
  • These methods are limited to handling discrete location spaces. The number of locations or cells is a crucial parameter that determines the perturbation mechanism and affects the privacy guarantee. Although it is possible to discretize continuous spaces, proper discretization is challenging to balance the trajectory utility and privacy.
  • Besides the discretization, methods based on the Exponential mechanism, such as NGram and ATP, suffer from high complexity in perturbation generation and trajectory utility loss when the discrete space is large. Additionally, designing public knowledge in NGram is challenging, especially for flying or sailing trajectories; synthesis methods in LDPTrace rely on the regularity of the training trajectories, which is not always guaranteed.
@@
Appendix B.1 details the weaknesses of the Exponential mechanism. In summary, it requires calculating the score functions of each possible pair of locations. This costs $\Theta(m)$ time, where $m$ is the number of locations in the discrete space. Additionally, each sampling (perturbation generation) also costs $\Theta(m)$ time. This time complexity is not affordable for large-scale location spaces.
@@
To address these limitations, we propose TraCS, a novel trajectory collection method that provides LDP guarantees for trajectories in continuous spaces.
3. Our Method: TraCS

### v2

@@
Abstract
Trajectory collection is fundamental for location-based services but often involves sensitive information, such as users’ daily activities, raising significant privacy concerns. Local Differential Privacy (LDP) provides strong privacy guarantees for users, even when the data collector is untrusted. Existing trajectory collection methods under LDP are limited to discrete location spaces, where the number of locations affects both privacy guarantees and trajectory utility. Moreover, many real-world scenarios, such as flying trajectories or sensor trajectories of wearable devices, operate in continuous location spaces, making existing methods inadequate.
@@ Abstract
This paper shifts the focus from discrete to continuous spaces for trajectory collection under LDP. We propose two novel methods: TraCS-D, which perturbs the direction and distance of locations, and TraCS-C, which perturbs the Cartesian coordinates of locations. Both methods are theoretically and experimentally analyzed for trajectory utility in continuous spaces. TraCS can also be applied to discrete spaces by rounding perturbed locations to the nearest discrete points. In this case, TraCS’s privacy and utility guarantees are independent of the number of locations in the space, and has only $\Theta(1)$ time complexity in each perturbation generation. Evaluation results on discrete location spaces validate the efficiency advantage and show that TraCS outperforms state-of-the-art methods with improved trajectory utility, especially for large privacy parameters.
1 Introduction
@@
Trajectory data from users—sequences of locations describing movement over time—are a fundamental resource for activity analysis and location-based services, including activity classification and routine detection. However, directly collecting trajectory data raises significant privacy concerns, especially given users’ growing privacy expectations and stringent regulations [1, 19]. Anonymization techniques like $k$-anonymity [28, 20] have proven insufficient to protect trajectory privacy [6]. Thus, stronger privacy guarantees are necessary for such data.
@@
Local differential privacy (LDP) provides a provable worst-case privacy guarantee by ensuring the hardness of distinguishing a user’s data from any released data. The privacy level of an LDP mechanism is quantified by its privacy parameter $\varepsilon$, which also generally implies the extent of data perturbation. Through LDP, users can release their perturbed trajectory data to an untrusted data collector while still benefiting from the services provided by the collector.
Existing LDP methods for trajectory collection are designed for discrete spaces. They leverage discrete LDP mechanisms, such as the Exponential mechanism [27], to perturb the trajectory data. Discrete LDP mechanisms are explicitly defined on the size of the location space. As a result, these methods either rely on properly splitting the entire location space into grids [29] or assume that the location space consists of a finite set of labeled locations (i.e. points of interest) [9, 37]. However, relying on discrete spaces has three issues: (i) Their privacy guarantees are inherently limited to the chosen discrete sets. For instance, a discrete location space with $10$ points always provides a weaker privacy guarantee than one with $100$ points, since any inference attack has at least a $10\%$ chance of success in the former, regardless of the privacy parameter $\varepsilon$. (ii) Their efficacy and efficiency are often constrained by the size of the discrete set. As the size increases, the probability of selecting the true location decreases. Meanwhile, the widely used Exponential mechanism incurs linear sampling complexity when generating each perturbed location, making it inefficient. (iii) Discrete methods are not directly applicable to naturally continuous location spaces, such as flying and sailing trajectories, sensor trajectories from wearable devices, and similar scenarios.
To address these limitations, this paper shifts the focus from discrete to continuous spaces for trajectory collection under LDP. A continuous space models locations as continuous values, such as GPS coordinates in $180,180]\times[-90,90]$. Unlike discrete spaces, it contains infinitely many candidate locations. Collecting trajectory data directly in continuous spaces is natural for many applications, and offers three key advantages over discretizing the space and then applying discrete methods: (i) The privacy guarantee holds for the entire continuous space, regardless of the “number” of locations. (ii) The sampling mechanism operates directly on the continuous space, so its efficacy and efficiency do not relate to the “number” of locations. (iii) Furthermore, perturbed locations in the continuous space can be post-processed by rounding to any discrete space within it, making the approach applicable to both continuous and discrete spaces.
Designing trajectory collection methods for continuous spaces under LDP presents unique challenges: they must ensure LDP for every location in the continuous space; they must preserve sufficient trajectory utility for downstream tasks compared to the original trajectories (formalized in Section 4); and they should efficiently generate perturbed locations to support real-time applications.
@@
This paper proposes two LDP methods for trajectory collection in continuous spaces. The main insight is to decompose the 22D continuous space into 11D subspaces and design mechanisms for each subspace. Our mechanisms leverage the concept of piecewise-based mechanisms [31, 21] for 11D bounded numerical domains. [^1] Another category of LDP mechanisms for bounded numerical domains is truncated mechanisms, e.g. the truncated Laplace mechanism [22, 15]. While they can be incorporated into TraCS, they are more complex and less effective than piecewise-based mechanisms. Section 3.5.4 provides details. Based on two different decompositions of the continuous space, we propose two methods: TraCS-D and TraCS-C. The method TraCS-D treats the continuous space as the composition of a direction space and a distance space. We exploit the direction information of trajectories and design a novel direction perturbation mechanism to ensure LDP for the direction space. Combined with a redesigned piecewise-based mechanism for the distance space, TraCS-D ensures LDP for the continuous 22D space. The method TraCS-C treats the continuous space as the Cartesian product of two distance spaces. Each location is represented by a 22D Cartesian coordinate, and TraCS-C perturbs these coordinates. Both methods guarantee LDP for the entire 22D continuous space. They can also be applied to any discrete space within the continuous space by rounding the perturbed locations to the nearest discrete points. Compared with existing methods based on the Exponential mechanism [9, 37], our methods are unaffected by the number of locations in the location space and require only $\Theta(1)$ time to generate each perturbed location. Specifically, our contributions are as follows:
@@
  • To the best of our knowledge, this is the first work to provide trajectory collection methods for continuous space under pure LDP. We propose TraCS-D and TraCS-C for this purpose. These methods introduce novel location perturbation mechanisms that exploit the direction information and coordinate information of trajectories in continuous spaces. We theoretically and experimentally analyze their trajectory utility.
@@
  • Our methods TraCS can also be applied to discrete spaces. Compared with existing methods for discrete spaces, TraCS has significantly lower computation complexity in generating perturbed locations. Experimental results on discrete location spaces validate the efficiency of TraCS, showing an average time cost of less than $1\%$ of NGram, L-SRR, and ATP [9, 37, 29], three state-of-the-art LDP methods designed for discrete spaces. Additionally, the results demonstrate that TraCS generally achieve better trajectory utility, particularly for larger privacy parameters.
Structure. The rest of the paper is organized as follows. Section 2 introduces the preliminaries. Section 3 presents TraCS-D and TraCS-C methods, their theoretical analysis, along with discussions on their differences and extensions. Following that, Section 4 presents the experimental evaluations on continuous and discrete spaces.
2 Preliminaries
@@
This section formulates the problem, introduces LDP, and reviews three trajectory collection methods under LDP, along with their shortcomings in continuous spaces, which motivates the necessity of designing new LDP methods for trajectory collection in continuous spaces.
2.1 Problem Formulation
@@
A typical trajectory collection schema consists of a set of trajectories from users and one collector. Each trajectory $\mathcal{T}$ is a sequence of locations $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$, where $\tau_{i}\in\mathcal{S}$ and $\mathcal{S}\subset\mathbb{R}^{2}$ is a continuous and bounded domain. The collector needs to collect these trajectories to provide analysis or location-based services.
@@
However, the collector is untrusted and may act as an adversary attempting to infer users’ sensitive data. Therefore, releasing the sensitive trajectories to the collector poses a privacy risk. To protect privacy, users perturb their trajectories using a privacy mechanism $\mathcal{M}:\mathcal{S}\to Range(\mathcal{M})$, and then send the perturbed trajectories $\mathcal{T}^{\prime}=\{\tau_{1}^{\prime},\tau_{2}^{\prime},\dots,\tau_{n}^{\prime}\}$ to the collector.
@@
We aim to design a mechanism $\mathcal{M}$ that satisfies $\varepsilon$-LDP, providing provable privacy guarantee for sensitive trajectories. Additionally, $\mathcal{M}$ should preserve utility, ensuring perturbed trajectories remain comparable to the originals, while incurring only modest computational overhead suitable for real-time applications.
2.2 Local Differential Privacy
Definition 1 ($\varepsilon$-LDP [11]).
@@ 2.2 Local Differential Privacy
A perturbation mechanism $\mathcal{M}:\mathcal{X}\to Range(\mathcal{M})$ satisfies $\varepsilon$-LDP, if for two arbitrary input $x_{1}$ and $x_{2}$, the probability ratio of outputting the same $y$ is bounded:
@@
If $\mathcal{M}(x)$ is continuous, the probability $\Pr[\cdot]$ is replaced by probability density function (pdfpdf). Intuitively, Definition 1 represents the difficulty of distinguishing $x_{1}$ and $x_{2}$ given $y$. Lower values of the privacy parameter $\varepsilon\in[0,+\infty)$ mean higher privacy. For example, $\varepsilon=0$ requires $\mathcal{M}$ to map two arbitrary inputs to any output $y$ with the same probability, thus the perturbed output contains no distribution information of the sensitive input, making any hypothesis-testing method to infer the sensitive $x$ powerless.
Theorem 1 (Sequential Composition of LDP [13, 37]).
@@
Sequential composition allows us to design LDP mechanisms for 22D data by composing two mechanisms for 11D.
2.2.1 Piecewise-based Mechanism
If $\mathcal{X}$ is continuous and bounded, the state-of-the-art LDP mechanisms are piecewise-based [31, 21], i.e. they sample the output from piecewise probability distributions defined on $\mathcal{X}$. We summarize them in the following form:
Definition 2 (Piecewise-based mechanism).
@@ 2.2.1 Piecewise-based Mechanism
where $p_{\varepsilon}$ is the sampling probability w.r.t. $\varepsilon$, $l_{x,\varepsilon},r_{x,\varepsilon})$ is the sampling interval w.r.t. $x$ and $\varepsilon$. $Range(\mathcal{M})\supseteq\mathcal{X}$ is also a continuous and bounded domain.
Piecewise-based mechanisms sample the output $y$ with a high probability $p_{\varepsilon}$ within $l_{x,\varepsilon},r_{x,\varepsilon})$ and a lower probability in the remaining two pieces, while satisfying the $\varepsilon$-LDP constraint. Typical instantiations include PM [31] for mean estimation and SW [21] for distribution estimation. Our perturbation mechanisms also leverage the piecewise-based design but with different instantiations tailored for different data domains.
2.2.2 $k$-$R$ and Exponential Mechanism
If $\mathcal{X}$ is discrete, especially when $|\mathcal{X}|$ is not large [32], a classical LDP mechanism is $k$-$R$ [18].
Definition 3.
$k$-$R$ is a sampling mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{X}$ that, given $|\mathcal{X}|=k$ and input $x\in\mathcal{X}$, outputs $y\in\mathcal{X}$ according to
@@ 2.2.2 $k$-$R$ and Exponential Mechanism
$k$-$R$ outputs the truth $x$ with a higher probability or outputs other values with a lower probability, while satisfying $\varepsilon$-LDP.
When there is a semantic distance (score) function between elements in $\mathcal{X}$, the Exponential mechanism [27] is more widely used. Unlike $k$-$R$, which treats other values (except $x$) equally, the Exponential mechanism leverages a score function to assign different probabilities. $k$-$R$ is a special case of the Exponential mechanism when the score function is a binary indicator function.
Definition 4 (Exponential Mechanism [27]).
Given a score function $d:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$, a privacy parameter $\varepsilon>0$, and a set of possible outputs $\mathcal{Y}$, the Exponential mechanism $\mathcal{M}:\mathcal{X}\to\mathcal{Y}$ is defined by:
$\Pr[\mathcal{M}(x)=y]=\frac{\exp\left(\frac{\varepsilon d(x,y)}{2\Delta d}\right)}{\sum_{y^{\prime}\in\mathcal{Y}}\exp\left(\frac{\varepsilon d(x,y^{\prime})}{2\Delta d}\right)},$
where $\Delta d=\max_{x,y,y^{\prime}\in\mathcal{Y}}|d(x,y)-d(x,y^{\prime})|$ is the sensitivity of the score function $d$.
[…]
2.3 Existing Methods
Table 1: Comparison with existing methods.
@@ 2.3 Existing Methods
NGram [9] Hierarchical decomposition Discrete space
L-SRR [29] Staircase $R$ mechanism
ATP [37] Direction perturbation
This paper (TraCS)
Direction* &
coordinate perturbation
Continuous space
@@
Table 1 summarizes three state-of-the-art trajectory collection methods under pure LDP and our methods.
NGram [9] hierarchically decomposes the physical space $\mathcal{S}$ into fine-grained discrete regions with semantics. Specifically, $\mathcal{S}$ is first decomposed into small regions $R_{s}$, which are further decomposed by categories into $R_{c}$, and then attached with time information $R_{t}$. A decomposed “gram” might look like $r_{sct}=\{\text{mountain, church, 3am}\}$. This decomposition allows trajectory $\mathcal{T}$ to be represented at a knowledge level, enabling the usage of public knowledge to define reachability, i.e. to constrain possible perturbed trajectories. For example, the above gram $r_{sct}$ is unlikely to be visited at 33am as a church is closed at that time, reducing the possible $\mathcal{T}^{\prime}$. Finally, NGram applies the Exponential mechanism to sample a perturbed trajectory from the reachable set.
L-SRR [29] introduces the staircase randomized response (SRR) mechanism for location perturbation. Unlike $k$-$R$, which assigns only two probabilities—one for the true location and another for all other locations—the SRR mechanism employs a staircase-shaped probability distribution. L-SRR recursively partitions the physical space $\mathcal{S}$ and forms location groups $G_{1},G_{2},\dots,G_{g}$ based on their distances to the true location. By integrating the SRR mechanism, L-SRR assigns higher probabilities to locations in groups that are closer to the true location, and lower probabilities to those farther away. This approach is similar to the Exponential mechanism, but uses group-level distance instead of location-level distance.
ATP [37] uses direction perturbation to limit the set of possible locations. It assumes that the location space is a finite set of labeled locations and represents a trajectory $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$ as a sequence of these locations. When perturbing a location $\tau_{i+1}$, ATP first perturbs the direction from $\tau_{i}$ to $\tau_{i+1}$. It divides the direction space into $k$ sectors and uses $k$-$R$ to sample a direction sector. Then, the perturbed location $\tau_{i+1}^{\prime}$ is constrained to be within the sampled sector, thereby limiting the possible $\tau_{i+1}^{\prime}$, which is then sampled using the Exponential mechanism.
@@
  • These methods are limited to handling discrete location spaces. The number of locations is a crucial parameter that determines the perturbation mechanism and affects the privacy guarantee. Although it is possible to discretize continuous spaces, proper discretization is challenging to balance the trajectory utility and privacy.
  • Beyond discretization challenges, approaches that rely on the Exponential mechanism (e.g. NGram and ATP) incur high computational cost to generate perturbations and tend to degrade trajectory utility. The other method L-SRR requires forming groups for every location in the space, which leads to the same scalability issue. Additionally, designing effective public knowledge (e.g. NGram) is challenging, particularly in continuous spaces where the location set is infinite and semantic information is difficult to define. Meanwhile, public knowledge is often dataset-specific and also known to the adversary, thereby undermining the privacy guarantee.
@@
Appendix C.1 details the weaknesses of the Exponential mechanism. In summary, before perturbing a location, it requires calculating the score functions $d(x,y)$ of all possible pairs of locations related to $x$. This costs $\Theta(m)$ time, where $m$ is the number of locations in the discrete space. Additionally, each sampling (generating each perturbed location) from the $m$-piece cumulative distribution function (CDF) of the Exponential mechanism also costs $\Theta(m)$ time. This time complexity is not affordable for large-scale location spaces. Moreover, since the sampling probability is determined by the score function, the spatial distribution and relative distances among locations further affect the probability assigned to each location, potentially resulting in utility loss.
@@
To address these limitations, we propose TraCS, a novel trajectory collection method that provides full LDP guarantees for trajectories in continuous spaces. TraCS is specifically designed to overcome the privacy, utility, and efficiency challenges posed by discretization.
3 Our Method: TraCS